# WakenLLM: Evaluating Reasoning Potential and Stability in LLMs via Fine-Grained Benchmarking

### v4

@@ 5 Conclusion
We present WakenLLM, a fine-grained framework for evaluating the Vague Perception phenomenon of LLMs. The framework supplies targeted cues that stimulate models to reflect on reasoning, substantially raising the empirical accuracy. Experiments show that while LLMs give up and answer Unknown on the first try, well-designed stimulation reliably steers them to the correct label and uncovers a persistent tendency toward conformity even after explicit reminding. These insights suggest training and evaluation protocols that value a model’s capacity for potential instead of the one-shot performance alone. Even though significant improvements in inner structures and parameter optimizations of LLMs may be limited in the near future, however, our findings suggest that current methods are still far from triggering LLMs’ full potential, and addressing this problem can bring unique prospective. Our research offers high underlying benefits regarding models’ potential.
@@ 6 Limitations
Even though our work has helped expose the empirical reasoning potential of LLMs, however, our final metrics are all calculated regarding labels, and outputting true labels with flawed reasoning process can be a common problem. How to fundamentally quantify their true ability with true reasoning process remains a problem in the future.
@@ 7 Ethical Considerations
Our research utilizes publicly available datasets (FLD, FOLIO, ScienceQA) containing synthetic logical reasoning problems without personal information. We acknowledge potential biases in our SFT and DPO training processes and implement atomic reasoning criteria to ensure fair evaluation. All training scripts, evaluation code, and prompt templates are made publicly available for reproducibility. We use LoRA fine-tuning to reduce computational requirements and focus on improving logical reasoning quality rather than enabling harmful applications. Our work adheres to academic integrity standards and contributes to responsible AI development through more interpretable reasoning.
Appendix A Baseline Details
@@
The CoT prompting method guides large language models (LLMs) to generate intermediate reasoning steps before producing a final answer. It improves multi-step reasoning performance by making the inference process explicit. This technique is particularly effective in arithmetic and logic-heavy tasks, as shown in prior works such as Wei et al. (2022b).
@@
Few-shot prompting involves presenting the model with a few annotated examples in the input to demonstrate the task format. It enables in-context learning without any gradient updates. Despite its simplicity, few-shot prompting serves as a strong baseline in many LLM applications (Brown et al., 2020).
@@
Tree-of-Thought extends the CoT framework by exploring multiple reasoning paths in a tree structure and evaluating intermediate steps. It mimics a planning process that considers alternative possibilities, significantly enhancing reasoning robustness and allowing backtracking or reevaluation of incorrect paths (Yao et al., 2023a).
@@
Plan-and-Solve separates the reasoning process into two stages: first planning a high-level outline or decomposition of the task, then solving sub-tasks based on that plan. This two-stage strategy enhances controllability and consistency in reasoning, and has shown strong performance across complex benchmarks (Wang et al., 2023a).
@@
Supervised Fine-Tuning (SFT) is the standard method for aligning LLMs by training on curated human-annotated question-answer pairs. Unlike prompt-only methods, SFT optimizes the model weights directly through gradient updates to produce desirable outputs. It forms the foundation of instruction-tuned models like InstructGPT and serves as a widely-used supervised baseline for alignment tasks (Ouyang et al., 2022). The common consensus is that SFT provides the essential starting point for alignment, but by itself is insufficient to fully capture complex human preferences, often requiring further preference optimization or reinforcement learning. The optimized formula is below:
@@
Direct Preference Optimization (DPO) is a reinforcement learning-free approach that fine-tunes LLMs using preference data. Given pairs of outputs (e.g., preferred vs. dispreferred), DPO directly optimizes the log-likelihood ratio between them, providing alignment benefits similar to reinforcement learning methods like PPO while avoiding their instability and complexity (Rafailov et al., 2023). The common consensus is that DPO offers a simpler and more stable alternative to reinforcement learning-based methods, making it a practical choice for preference alignment, though it may still struggle with capturing nuanced or long-term user preferences. The formula can be concluded as below:
@@
A.1 Experiment Setup
We choose four LLMs: GPT-4o, GPT-3.5-turbo-1105, LLaMA-3.1-8B and Qwen-2.5-7B for our detailed baseline experiments. We categorize our baseline methods into two kinds: training and non-training, in order to see whether the Vague Perception phenomenon can be solved with or without parameter revision. For model training and open-source model deployment, the experiments ran on four Tesla V100-SXM2-32GB GPUs.
A.2 Training Methods Details
As explained before in Appendix A.1, we elaborate our facilities and setup for training. In this section, we demonstrate our details for the training.
For SFT (Supervised Fine-Tuning), we extract 1,000 training samples from exactly same dataset, the prompt instruction can be found in Appendix D.2.
For DPO (Direct Preference Optimization), we apply LLM Judge to help select the reasoning process with many atomic steps as possible. This includes two parts: (1) Prompt Template that includes task guiding details (e.g., definition of atom steps) and task instructions (2) Testing cases that can perfectly prove and verify which response is better given the metrics and requirements. All information above can be found in Appendix D.2.
Appendix B Benchmark Details
FLD
Following Morishita et al. (2024), FLD is a fully synthetic benchmark for multi‑step logical inference. Formal proof chains of depth 0–15 are automatically generated and then verbalized into natural‑language premises plus a hypothesis. Each instance is labeled Proved, Disproved, or Unknown, making the dataset ideal for assessing whether a model can (i) draw correct conclusions and (ii) abstain when evidence is insufficient.
FOLIO
FOLIO (Han et al., 2022) contains 1,430 human‑written first‑order‑logic stories that bridge everyday commonsense with formal theorem proving. Every narrative provides its gold logical forms and a hypothesis whose truth value (True, False, or Unknown) is automatically certified. The official split is 1 001/226/203 for train/dev/test, offering fewer but richer problems than FLD and requiring explicit refusals when the passage leaves the hypothesis undecided.
ScienceQA
ScienceQA (Lu et al., 2022) is a multi-modal multiple‑choice benchmark with 21,208 elementary‑ and middle‑school questions across 26 topics. To create an abstention setting comparable to the logic corpora, we select questions from Physics and Biology (science) as well as Writing‑Strategies and Figurative‑Language (arts). For 50% of the chosen items we delete two random supporting sentences with GPT‑4o, then manually verify that no unique answer remains; these items are re-labeled UNKNOWN. The final corpus therefore mixes standard A–D supervision with an explicit refusal class.
Topic Analysis
Unknown Labeling
For the first FLD and FOLIO, the Unknown label is originally included in their annotations. For ScienceQA, we randomly select half of the samples and employ GPT-4o to randomly delete two sentences from the given facts. We then re-label these modified samples as Unknown. Additionally, we perform manual verification to ensure that each relabeled sample is semantically consistent with its assigned label.
Ethical Statement
All datasets are distributed for non‑commercial research under permissive licenses (FLD: MIT; FOLIO: research‑only; ScienceQA: CC‑BY‑NC‑SA 4.0). Our sentence‑removal augmentation introduces no personal data and was human‑verified to avoid unsafe or misleading content.
Figure 6: Scope of Vague Perception. This phenomenon underlies a wide range of reasoning failures. By identifying and sub-categorizing cases that are potentially solvable, it reveals a substantial portion of failures that stem from Vague Perception rather than fundamental incapacity. Addressing this issue can significantly advance fine-grained reasoning and improve the overall trustworthiness of LLMs.

### v5

@@ 6 Conclusion
In this work, we formulate the Vague Perception phenomenon: the inclinations of LLMs outputting unknown despite being potentially able to solve input questions, and such abstention behavior strongly influence LLM performances. We present WakenLLM, a fine-grained pipeline for evaluating and mitigating the Vague Perception phenomenon in LLM reasoning. Our pipeline elicits LLMs’ potential reasoning ability, substantially improving the accuracy among tested datasets. Experiments show that while LLMs give up and answer unknown on the first try, a well-designed pipeline can stimulate them to answer correctly. These insights suggest evaluation metrics that value LLMs’ ability beyond one-shot performances alone. By studying Vague Perception, Conformity and Degradation, we better understand reasoning behaviors vary from model families and parameter sizes. Although significant leaps in LLM architectures may be limited in the near future, our findings suggest that current reasoning baselines remain far from eliciting LLMs’ full potential, paving the way for future perception-aware LLM designing. Our research offers high values for trustworthiness and reliable LLM reasoning development.
@@ 7 Limitations
Even though our work has helped elicit LLMs’ reasoning potential, our final metrics are calculated with respect to output labels, and producing the correct label while providing a flawed reasoning chain is common and remains an open problem for better reinforcement learning approaches in the future. How to fundamentally quantify their true ability with true reasoning process remains a problem in the future. Besides, specifying whether the output Unknown is due to understanding the unverifiable elements or inability to fully comprehend still remains a problem, currently we just clarify demands within prompts.
@@ 8 Ethical Considerations
Our research utilizes publicly available datasets (FLD, FOLIO and ScienceQA) containing synthetic and artificial reasoning samples without personal information. We acknowledge potential biases in our SFT and DPO training processes and implement atomic reasoning criteria to ensure fair evaluation. All training scripts, evaluation code, and prompt templates are made available for reproducibility. We focus on improving reasoning quality rather than enabling harmful applications. Our work adheres to academic integrity standards and contributes to responsible AI development through more interpretable reasoning.
Appendix A Related Work Details
Chain-of-Thought (CoT) prompting (wei2022cot) first showed that furnishing step-wise exemplars markedly improves LLM performance on arithmetic and commonsense problems. Follow-up work refines this paradigm: scratchpads (nye2021scratchpad), least-to-most prompting (zhou2022ltm), and self-consistency (wang2023selfconsistency). More recently, multi-agent and symbolic reasoning approaches like multiagent debate (du2024multiagentdebate), symbolic chain-of-thought (xu2024symboliccot), and neurosymbolic reasoning with logic provers (olausson2023linc) further strengthen factuality and logical consistency.
Tool-Augmented Reasoning
Program-Aided LMs (gao2022pal) delegate arithmetic or logical computation to Python. ReAct (yao2022react) interleaves reasoning and external tools. CRITIC (gou2024critic) validates intermediate answers with execution tools. Chain-of-Verification (dhuliawala2024chainofverification) reduces hallucinations by verifying sub-claims against evidence.
Iterative Self-Refinement and Critique
Self-Refine (madaan2023selfrefine) establishes a generate–criticise–rewrite loop. Ambiguity handling and uncertainty decomposition are also explored: AmbigQA (min2020ambigqa), Decomposing Uncertainty (hou2024decomposeuncertainty), and rank-based calibration (huang2024rankcalibration).
Datasets for Reasoning and Verification
Classic benchmarks like BoolQ (clark2019boolq), SQuAD 2.0 (rajpurkar2018squad2), FEVER (thorne2018fever), and HotpotQA (yang2018hotpotqa) stress-test fact verification, ambiguity, and multi-hop reasoning. Meanwhile, multilingual and cross-domain robustness is increasingly crucial for real-world deployment. Recent work such as MuLA-F (li-etal-2025-multilingual-federated) highlights the importance of handling multilingual heterogeneity and collaborative anomaly detection across diverse social media environments.
Active Prompting
Active-Prompt (diao2024activeprompting) selects informative exemplars to boost performance under budget constraints. Meanwhile, ling2025instructionboundaryquantifyingbiases denotes different prompt coverage leads to various effects.
Faithfulness and Safety
Beyond correctness, methods ensure interpretability and reliability, such as Faithful-CoT (lyu2023faithful), symbolic verification (xu2024symboliccot), and multi-agent debate (du2024multiagentdebate). Recent research on trustworthy language models has focused on two major directions: mitigating misinformation and developing watermarking techniques for tracing generated content. Liu et al. liu2025surveyproactivedefensestrategies survey proactive defenses against misinformation in LLMs, proposing a three-pillar framework for improving model reliability. In related representation learning work, Liu et al. liu-etal-2022-hiure introduce a hierarchical contrastive approach for unsupervised relation extraction, offering insights into structured semantic modeling. A parallel line of studies explores watermarking for text and multimodal generation. Zhang et al. zhang2025cohemarknovelsentencelevelwatermark propose COHEMARK, a sentence-level watermark that preserves semantic coherence. Liu et al. liu2025vlamarkcrossmodalwatermark extend watermarking to vision–language models, while Huo et al. huo-etal-2025-pmark introduce a distortion-free semantic watermark with channel constraints. More recently, Zhang et al. zhang-etal-2025-catmark present CATMARK, a context-aware thresholding framework for robust cross-task watermarking.
Appendix B Baseline Details
In this section, we run multiple reasoning frameworks that help improve downstream tasks of LLM reasoning. By comparing our proposed pipeline with existing baselines, we can further prove that current frameworks are far from eliciting LLMs’ full potential. Detailed templates and examples can be found in Appendix N.
@@ Appendix B Baseline Details
The CoT prompting method guides large language models (LLMs) to generate intermediate reasoning steps before producing a final answer. It improves multi-step reasoning performance by making the inference process explicit. This technique is particularly effective in arithmetic and logic-heavy tasks, as shown in prior works such as wei2022chain.
@@
Few-shot prompting involves presenting the model with a few annotated examples in the input to demonstrate the task format. It enables in-context learning without any gradient updates. Despite its simplicity, few-shot prompting serves as a strong baseline in many LLM applications (brown2020language).
@@
Tree-of-Thought extends the CoT framework by exploring multiple reasoning paths in a tree structure and evaluating intermediate steps. It mimics a planning process that considers alternative possibilities, significantly enhancing reasoning robustness and allowing backtracking or reevaluation of incorrect paths (yao2023tree).
@@
Plan-and-Solve separates the reasoning process into two stages: first planning a high-level outline or decomposition of the task, then solving sub-tasks based on that plan. This two-stage strategy enhances controllability a in reasoning, and has shown strong performance across complex benchmarks (wang2023plan).
@@
Supervised Fine-Tuning (SFT) is the standard method for aligning LLMs by training on curated human-annotated question-answer pairs. Unlike prompt-only methods, SFT optimizes the model weights directly through gradient updates to produce desirable outputs. It forms the foundation of instruction-tuned models like InstructGPT and serves as a widely-used supervised baseline for alignment tasks (ouyang2022training). The common consensus is that SFT provides the essential starting point for alignment, but by itself is insufficient to fully capture complex human preferences, often requiring further preference optimization or reinforcement learning. The optimized formula is below:
@@
Direct Preference Optimization (DPO) is a reinforcement learning-free approach that fine-tunes LLMs using preference data. Given pairs of outputs (e.g., preferred vs. dispreferred), DPO directly optimizes the log-likelihood ratio between them, providing alignment benefits similar to reinforcement learning methods like PPO while avoiding their instability and complexity (rafailov2023direct). The common consensus is that DPO offers a simpler and more stable alternative to reinforcement learning-based methods, making it a practical choice for preference alignment, though it may still struggle with capturing nuanced or long-term user preferences. The formula can be concluded as below:
@@
Self-Reflection
Self-reflection has been widely adopted as a strategy for improving LLM reasoning by prompting the model to critique and refine its own initial answer. However, this paradigm implicitly assumes that the model can (i) detect flaws in its original reasoning, (ii) identify missing or contradictory evidence, and (iii) reliably revise its conclusion. Prior work shows that these assumptions rarely hold in practice: when the initial reasoning is misguided, the reflection stage often reproduces the same errors, leading to self-confirming loops rather than genuine correction. In addition, self-reflection primarily operates on the model’s final answer, without addressing perception-level uncertainty or the model’s tendency to prematurely output Unknown.
@@
B.1 Experiment Setup Details
We choose four LLMs: GPT-4o, GPT-3.5, LLaMA-3.1-8B and Qwen-2.5-7B for our detailed baseline experiments. We categorize our baseline methods into two kinds: training and non-training, in order to see whether the Vague Perception phenomenon can be solved with or without parameter revision. For model training and open-source model deployment, the experiments ran on four Tesla V100-SXM2-32GB GPUs.
B.2 Training Methods Details
As explained before in Appendix B.1, we elaborate our facilities and setup for training. In this section, we demonstrate our details for the training.
For SFT (Supervised Fine-Tuning), we extract 1,000 training samples from exactly same dataset, the prompt instruction can be found in Appendix N.
For DPO (Direct Preference Optimization), we apply LLM Judge to help select the reasoning process with many atomic steps as possible. This includes two parts: (1) Prompt Template that includes task guiding details (e.g., definition of atom steps) and task instructions (2) Testing cases that can perfectly prove and verify which response is better given the metrics and requirements. All information above can be found in Appendix N.
Table 7: Detailed listing of all metrics used among different experimental settings in our work. Metrics within different settings or systems are separated by dashed lines. RP stands for reasoning process included within the setting. $f\in\{s,u\}$ indicates whether the sample types belong to solvable or unsolvable. $n\in\{1,2\}$ indicates the stage of processing. Omitting subscripts indicates all sample types; omitting superscripts indicates all two stage considered. Detailed calculations can be found in Appendix F.
Vanilla Settings Definition Rationality/Motivation
True Converting (TC) $TC^{n}_{t}$ represents samples converted to correct labels. Measuring LLMs’ true potential under stimulation.
True Converting Rate (TCR) $TCR^{n}_{t}$ is the percentage of $TC^{n}_{t}$. Calculate the percentage of LLMs’ potential of solving problems after stimulation.
False Converting (FC) $FC^{n}_{t}$ represents samples converted to incorrect labels. Under stimulation, problems LLMs are stimulated to produce wrong answers instead of giving up.
False Converting Rate (FCR) $FCR^{n}_{t}$ is the percentage of $FC^{n}_{t}$. Research the inclination of LLMs outputting wrong answer over abstention (answering incorrectly).
Unknown Converting (UC) $UC^{n}_{t}$ represents samples converted to unknown samples. To confirm that the samples are too hard for LLMs to reason.
Unknown Converting Rate (UCR) $UCR^{n}_{t}$ is the percentage of $UC^{n}_{t}$. Research the inclination of LLMs abstention (answering unknown) over outputting wrong answer.
Overall Converting Rate (OCR) After two stage stimulation, the percentage of Vague Perception samples are converted to TC, sum of $TCR^{1}$ and $TCR^{2}$. Potential part of samples LLMs can reason correctly, but did wrong because of abstention.
Remind-then Guide (Label&RP) Definition Rationality/Motivation
Label Conformity LLMs exhibit distinct tendency of conformity on solvable and unsolvable labels, denoted as $Conf_{t}$. Research how conformity biases are enlarged facing different label forms.
Correct Guiding Rate (CGR) Taking the previous reasoning trace as input, to determine whether performance is better or worse with reflection. LLM training based on RL is result-oriented, leaving the reasoning trace flawed and biased.
Reasoning Process Coherence (RPC) Based on original output, LLMs should rethink the problem, this is to define whether the reasoning trace is consistent. As the stability of reasoning is important, ensuring the consistency of reasoning can be important.
Other Experiments Definition Rationality/Motivation
Degradation (Deg) In stage 2 stimulation, the input $FC^{2}$ may be degraded and output unknown, which should be denoted as $UC^{2}$. Consideration for randomness and instability of LLMs output. Better make up for the pipeline.
Appendix C Datasets Details
FLD
Following morishita2024enhancing, FLD is a fully synthetic benchmark for multi‑step logical inference. Formal proof chains of depth 0–15 are automatically generated and then verbalized into natural‑language premises plus a hypothesis. Each instance is labeled Proved, Disproved, or Unknown, making the dataset ideal for assessing whether a model can (i) draw correct conclusions and (ii) abstain when evidence is insufficient.
FOLIO
FOLIO (han2022folio) contains 1,430 human‑written first‑order‑logic stories that bridge everyday commonsense with formal theorem proving. Every narrative provides its gold logical forms and a hypothesis whose truth value (True, False, or Unknown) is automatically certified. The official split is 1 001/226/203 for train/dev/test, offering fewer but richer problems than FLD and requiring explicit refusals when the passage leaves the hypothesis undecided.
ScienceQA
ScienceQA (lu2022learn) is a multi-modal multiple‑choice benchmark with 21,208 elementary‑ and middle‑school questions across 26 topics. To create an abstention setting comparable to the logic corpora, we select questions from Physics and Biology (science) as well as Writing‑Strategies and Figurative‑Language (arts). For 50% of the chosen items we delete two random supporting sentences with GPT‑4o, then manually verify that no unique answer remains; these items are re-labeled Unknown.
Fields/Domains Analysis
Figure 6: Leaderboard ranking (left) and average experimental results across datasets (right) in Figure 3.
Unknown Label Annotation
For the first FLD and FOLIO, the Unknown label is originally included in their annotations. For ScienceQA, we randomly select half of the samples and employ GPT-4o to randomly delete two sentences from the given facts. We then re-label these modified samples as Unknown. Additionally, we perform manual verification to ensure that relabeled samples are semantically consistent with assigned labels, which are objectively unverifiable.
Search-Based Planning Frameworks
Instead of generating a single chain, Tree-of-Thought (ToT) views partial solutions as nodes in a tree and performs breadth-first search with value heuristics (yao2023tot). Decomposed Prompting turns a complex task into callable sub-prompts that can be solved independently (khot2023decomposedpromptingmodularapproach).
Appendix D Metrics Details
As shown in Table 7, we elaborate all used metrics in our experiment, along with the motivation of designing them. A solid and comprehensive evaluation system offers contributions to LLM reasoning.
Appendix E LLM Usage Statement
GPT-4o, Claude-3-7 and Grok 3 are used for code organizing and paper framing.
Appendix F Formula Details
In the main body of paper, we highly focus on True Converting (TC) flow that LLMs are able to convert Vague Perception samples to predict correct labels. In this section we explain False Converting (FC) and Unknown Converting (UC) in detail. As we define the whole evaluation system, we also map out their relationships.
F.1 False and Unknown Converting
As demonstrated in 3.2, we denote the notion of TCR. Likewise, we define the concepts regarding false converting and unknown converting similarly:
$FCR_{t}^{n}\;=\;\frac{\bigl|FC_{t}^{n}\bigr|}{\bigl|D_{\mathrm{VP}}^{n}\bigr|},\quad UCR_{t}^{n}\;=\;\frac{\bigl|UC_{t}^{n}\bigr|}{\bigl|D_{\mathrm{VP}}^{n}\bigr|}$
Note that, Unknown Converting (UC) is still categorized into Vague Perception phenomenon: Outputting unknown when facing solvable samples. We can tell the following relationship for any stage that n can be 1 or 2:
$UCR_{t}^{n}+FCR_{t}^{n}+TCR_{t}^{n}\;=\;\mathbbm{1}$
$UC_{t}^{n}\cup FC_{t}^{n}\cup TC_{t}^{n}\;=\;D_{VP}^{n}$
$UC_{t}^{n}\cap FC_{t}^{n}\cap TC_{t}^{n}\;=\;\emptyset$
F.2 Overall Converting & Latent Ability
We clarify the relationship between Overall Converting Rate (OCR) in Section 3.2 with the latent ability in 5.5, as we depicted, by comparing True Converting (TC) with $D_{VP}$, we can get the ability of how many samples can be eventually converted to right labels within all Vague Perception samples, whereas the latter indicates overall accuracy gain among the all samples D in a broader aspect.
F.3 Remind-then-Guide Label Conformity
In this section, we talk about the formula of conformity in detail. We define one single-turn conformity as below:
$Conf_{t}^{n}=TCR^{n}_{t}-TCR_{t}^{n^{\prime}}$
During the experiments, we left out the influence of stage-wise performances, and focus on the true converting rate (TCR) changes separately on solvable and unsolvable samples, acquired with and without assigned labels. As the conformity is always considered to be a model attribute rather than a random phenomenon, we talk about the conformity regarding sample types, we pay less attention to the stage-wise analysis. By averaging the two-stage conformity, the formula in our experiments can be decomposed as:
$Conf_{t}$ $=\frac{1}{2}(TCR^{1}_{t}-TCR_{t}^{1^{\prime}})+\frac{1}{2}(TCR^{2}_{t}-TCR_{t}^{2^{\prime}})$
$=\frac{1}{2}\frac{|TC^{1}_{t}|-|TC_{t}^{1\prime}|}{|D^{1}_{VP}|}+\frac{1}{2}\frac{|TC^{2}_{t}|-|TC_{t}^{2\prime}|}{|D^{2}_{VP}|}$
Appendix G Leaderboard Analysis
As illustrated in Figure 6, the conformity towards different label formats vary considerably across models. Moreover, the quality of reasoning traces does not necessarily correlate with the final accuracy. This observation implies that LLMs can often produce correct answers despite flawed or shallow reasoning processes. We hypothesize that this phenomenon stems from the reinforcement learning (RL)-based training paradigm, which primarily optimizes for outcome-level rewards rather than reasoning quality. As a result, models tend to acquire “good answers with bad reasoning”—a form of outcome–process misalignment that contradicts the long-standing cognitive principle that good thinking leads to good answers (turpin2023language; wang2024reinforced; ziegler2019fine; christiano2017deep).
Parameters Analysis
According to the results, we can tell that except for GPT-3.5 openai2022gpt35 being an outdated model, for Qwen qwen2024qwen25 and LLaMA meta2025llama3 series, models with smaller parameters sizes tend to perform better. The takeaway aligns well with our findings in Figure 5 and Section 5.7, as higher the parameters, the more likely LLMs are to be cautious, whereas models with smaller parameters tend to be more instable and influenced by the experimental settings, this trade-off can be of guidance for models selection for different requirements.
Appendix H Detailed RtG Conformity Analysis
In this section, we dive deeper into the conformity effects under four different misguiding rate (M=0%, 50%, 66.67% and 100%): the percentage of input labels are replaced with the wrong ones. Detailed results are shown in Table 91011. Generally, all models demonstrate stronger conformity effects regarding higher misguiding rate.
Examining Table 9 reveals that misguidance has a pronounced influence on both stages of converting. For GPT-3.5, the FLD dataset exhibits a $TCR^{1}$ of only $5.83\%$ at a 100% misguiding rate, but this climbs to $54.17\%$ when M drops to zero. Similar trends appear on FOLIO and Arts, where reducing $M$ from $100\%$ to $0\%$ increases $TCR^{1}$ from $2.70\%$ to $18.92\%$ and from $22.92\%$ to $37.50\%$ respectively. GPT-4o exhibits slightly higher tolerance to misguidance: on FLD its $TCR^{1}$ rises from $15.27\%$ at $M=100\%$ to $26.22\%$ at $M=0\%$, and on Science from $8.42\%$ to $20.26\%$. However, even for GPT-4o, increasing misguidance still reduces True Converting Rate (TCR) largely, illustrating the general monotonic relationship between misguiding rate and conformity.
The RtG analysis for LLaMA-3.1 series models in Table 10 shows a striking gap in robustness. On the FLD dataset, the 8B model’s $TCR^{1}$ climbs from $0.36\%$ at $M=100\%$ to $75.18\%$ when there is no misguidance, whereas the 405B variant rises only from $1.90\%$ to $17.66\%$ Similar improvements for the 8B model across FOLIO and Science confirm that smaller models can recover significantly more reasoning potential when labels are correctly provided. By contrast, the 405B model shows modest gains and remains low on Arts, suggesting that larger models are more vulnerable to misguidance.
Table 11 compares Qwen-2.5-7B and Qwen-2.5-72B. For FLD and FOLIO, the 72B model performances better with less misguidance than the 7B model, with $TCR^{1}$ increasing from $13.25\%$ to $77.71\%$ on FLD and from $19.18\%$ to $42.47\%$ on FOLIO han2022folio. Yet on Science and Arts, the 7B model attains higher $TCR^{1}$ at $M=0\%$, reaching $50.83\%$ on Science and $55.93\%$ on Arts versus $37.84\%$ and $24.22\%$ for the 72B variant. These observations highlight that RtG counts on not only on models size but also on dataset fields.
0: Dataset $\mathcal{D}=\{(x_{i},y_{i})\}_{i=1}^{N}$ with labels $y_{i}\in\{\mathrm{T},\mathrm{F},\mathrm{U}\}$;base LLM model $\mathcal{M}$; two stimulation prompts $P_{1}$ and $P_{2}$
0: Stage-1 categories $\mathrm{TC}^{1},\mathrm{UC}^{1},\mathrm{FC}^{1}$ and Stage-2 categories $\mathrm{TC}^{2},\mathrm{FC}^{2}$
[…]
2:for all $(x_{i},y_{i})$ in $\mathcal{D}$ do
3:  $\hat{y}_{i}\leftarrow\mathcal{M}(x_{i})$ // first-pass prediction
4:  if ( $y_{i}\in\{\mathrm{T},\mathrm{F}\}$ and $\hat{y}_{i}=\mathrm{U}$ ) or ( $y_{i}=\mathrm{U}$ and $\hat{y}_{i}=\mathrm{U}$ and model lacks reasoning ) then
5:   $\mathrm{D_{VP}}\leftarrow\mathrm{D_{VP}}\cup\{(x_{i},y_{i})\}$ // collect Vague Perception samples
6:  end if
7:end for
[…]
9:for all $(x_{i},y_{i})$ in $\mathrm{D_{VP}}$ do
10:  $\hat{y}_{i}^{(1)}\leftarrow\mathcal{M}(\text{stimulation}(x_{i},P_{1}))$ // first-stage stimulation
11:  if $\hat{y}_{i}^{(1)}=y_{i}$ then
12:   $\mathrm{TC}^{1}\leftarrow\mathrm{TC}^{1}\cup\{(x_{i},y_{i})\}$ // True Converting at stage 1
13:  else if $\hat{y}_{i}^{(1)}=\mathrm{U}$ then
14:   $\mathrm{UC}^{1}\leftarrow\mathrm{UC}^{1}\cup\{(x_{i},y_{i})\}$ // Unknown Converting at stage 1
15:  else
16:   $\mathrm{FC}^{1}\leftarrow\mathrm{FC}^{1}\cup\{(x_{i},y_{i})\}$ // False Converting at stage 1
17:  end if
18:end for
[…]
20:for all $(x_{i},y_{i})$ in $\mathrm{FC}^{(1)}$ do
21:  $\hat{y}_{i}^{2}\leftarrow\mathcal{M}(\text{stimulation}(x_{i},P_{2}))$ // second-stage stimulation
22:  if $\hat{y}_{i}^{2}=y_{i}$ then
23:   $\mathrm{TC}^{2}\leftarrow\mathrm{TC}^{2}\cup\{(x_{i},y_{i})\}$ // Corrected at stage 2
24:  else
25:   $\mathrm{FC}^{2}\leftarrow\mathrm{FC}^{2}\cup\{(x_{i},y_{i})\}$ // Still incorrect at stage 2
26:  end if
27:end for
28:return $\mathrm{TC}^{1},\mathrm{UC}^{1},\mathrm{FC}^{1},\mathrm{TC}^{2},\mathrm{FC}^{2}$
Algorithm 1 Vanilla Setting Pipeline of the WakenLLM Pipeline.